Type a random 6-bits binary number, then converted to decimal. Show the steps.
Answer #1

Answer:

So lets start with typing a 6-bits binary number

take it as 101010

so lets start

0 * 2^0 = 0 * 1= 0

1 * 2^1 = 1 * 2 = 2

0 * 2^2 = 0 * 4 = 0

1 * 2^3 = 1* 8 = 8

0 * 2^4 = 0*16 = 0

1 * 2^5 = 1* 32 = 32

So basically what we have to do is start multiplyng each number from left with 2^n where n=0,and further going to right increasing it by 1.

so in this example starting from left

0*(2^0)=0

and further going right 1*(2^1)=2,here n increases by 1,n=2.

and further going on.

now add up all the products

like in this exampe

0+2+0+8+0+32=42 which is the answer.

Please provide a thumbs up feedback.It encourages us.Have a Good day):.